OpenStudy (kittiwitti1):

A rubber ball weighs 49 N. a) What is the mass of the ball? (Answer: 5.0 kg) b) What is the acceleration of the ball if an upward force of 69 N is applied? (Answer: 4.0 m/s/s) Medal for correct solving process? :) First answer shows my work.

OpenStudy (kittiwitti1):

For a) I used the equation W = Fg = mg, which becomes 49 = m(9.8). This approximates m to 5.0 kg. --- For b) I used the equation \[\Sigma Fy = Fa - W = ma\] which becomes 69-49 = 20 N. I then use equation F = ma which becomes 20 = 5.0a, which makes a \[4.0 m/s ^{2}\]